# ProSAM: Enhancing the Robustness of SAM-based Visual Reference Segmentation with Probabilistic Prompts

### v2

@@ Abstract
The recent advancements in large foundation models have driven the success of open-set image segmentation, a task focused on segmenting objects beyond predefined categories. Among various prompt types (such as points, boxes, texts, and visual references), visual reference segmentation stands out for its unique flexibility and strong zero-shot capabilities. Recently, several SAM-based methods have made notable progress in this task by automatically generating prompts to guide SAM. However, these methods often generate prompts at object boundaries due to suboptimal prompt encoder, which results in instability and reduced robustness. In this work, we introduce ProSAM, a simple but effective method to address the stability challenges we identified in existing SAM-based visual reference segmentation approaches. By learning a variational prompt encoder to predict multivariate prompt distributions, ProSAM avoids generating prompts that lie in unstable regions, overcoming the instability caused by less robust prompts. Our approach consistently surpasses state-of-the-art methods on the Pascal-5i and COCO-20i datasets, providing a more robust solution for visual reference segmentation.
@@ 1 Introduction
Figure 1: The same mask can be generated by SAM using various prompts. Specifically, different colors represent distinct object regions in the latent prompt embedding space, with each region corresponding to a unique object in the image. Arrows represent the one-to-one mapping between latent prompt embeddings and 2D coordinates within the image.
To tackle the challenges resulting from long-tailed data shortage and descriptive limitation, image segmentation with visual references (i.e., an annotated reference image that indicates the objects of interest) has become increasingly important. Its ultimate goal is to segment similar objects as indicated in the annotated reference image, regardless of the semantic category of target objects. Recently, various methods [45, 20, 32] have leveraged the exceptional segmentation capabilities of SAM, achieving significant breakthroughs in visual reference segmentation. The main idea of these methods is to generate prompts that direct SAM to predict masks for the target objects. Among these approaches, training-based methods [32] that learn prompt embeddings to guide SAM mask generation have achieved state-of-the-art (SOTA) performance. However, the existing training approaches often direct the prompts toward the boundaries between different objects, resulting in instability and reduced robustness. This robustness challenge is specifically caused by the inherent design of SAM — identical masks can be generated from different prompts (see Figure 1). This oversight poses a significant challenge for the training-based methods, where stability and generalizability are essential for zero-shot capability.
@@
  • We identify a commonly overlooked limitation in the SOTA SAM-based visual reference segmentation approach, where prompts are often generated at object boundaries, leading to instability and reduced robustness.
@@ 3.2 Existing Training-Based Method Using SAM
In essence, VRP-SAM focuses solely on the mask-level differences, while overlooking the potential for further optimizing the prompt encoder to generate more robust prompts.
@@ 4.1 Robustness Challenge
Figure 3: The intuitive illustration of the high-level idea behind the proposed variational prompt encoder. The dashed arrow indicates the sampling and prompting procedure during the training phase, while the solid arrow represents the prompting strategy employed during inference.
@@ 4.2 Variational Prompt Encoder
A straightforward instantiation of $P(\bm{z}|I_{r},M_{r},I_{t})$ is assuming $\bm{z}$ follows a conditional multivariate Gaussian distribution. However, considering our motivation is to push the mean prompt $\bm{\hat{\mu}}_{\bm{z}}$ toward the center of the target prompt embedding region, a more heavy-tailed t-distribution would be a better choice due to its statistical properties of having more chance to sample outliers [5]. In other words, when $\bm{\hat{\mu}}_{\bm{z}}$ is close to the boundary of $\mathcal{R}_{I_{r},M_{r},I_{t}}$, it is more likely to sample a prompt falling outside the target prompt embedding region (green region in Figure 3) if it follows heavy-tailed t-distribution. Given that the outlier prompt is more likely to result in low-quality mask predictions, the force to push the mean prompt toward the center of $\mathcal{R}_{I_{r},M_{r},I_{t}}$ will be greater with larger gradients. Inspired by Kim et al. [12], we formulate our variational prompt distribution $q_{\phi}(\bm{z}|I_{r},M_{r},I_{t})$ with the heavy-tailed property and diagonal covariance as follows,
@@ 5.1 Experimental Setup
Methods Image Encoder Learnable Params COCO-20i PASCAL-5i
@@ F-0 F-1 F-2 F-3 Mean F-0 F-1 F-2 F-3 Mean
Implementation Details. To ensure a strictly fair comparison with VRP-SAM, the SOTA method in the visual reference segmentation, we ensure that all the experimental settings (e.g., random seed, LR scheduler, optimizer) and hyper-parameters (e.g., number of layers, prompt embedding dimensions, number of prompts) are identical to VRP-SAM. Specifically, the model architecture of the variational prompt encoder is similar to VRP-SAM (see section 8 in Appendix for more details), except two linear layers have been added at the end of the prompt encoder to predict the mean and log standard deviation of the multivariate prompt distribution. In other words, our method is easy to implement with only a few lines of code, but can effectively boost the robustness of generated prompts. During training, the number of Monte Carlo samples K has been set to 10, with the degrees of freedom $\nu$ of 5. Also, we employ the AdamW [21] optimizer with weight decay of 1e-6 and the cosine annealing learning rate scheduler with warm restart [22] after 15 epochs, and the initial learning rate of 1e-4. The model is trained with 100 epochs with a fixed random seed of 321. For the choice of image encoder utilized by the variational prompt encoder, ResNet-50 [7] is adopted following VRP-SAM. In addition, our variational prompt encoder predicts 50 multivariate prompt distributions per target image and only uses 50 mean prompts during inference, which guarantees the same number of prompts have been used in the experimental study for both VRP-SAM and ProSAM. In terms of the visual annotation types, we follow the same procedure as VRP-SAM [32] and SEEM [46] to automatically generate points, scribbles, and boxes based on the mask annotations. Lastly, the mean intersection over union (mIOU) is adopted to evaluate our segmentation performance across all datasets. All experiments on COCO-20i were conducted on 4 RTX 4090 GPUs with a batch size of 2 per GPU, whereas the experiments on Pascal-5i were conducted on 1 H100 GPU with a batch size of 2.
@@ 5.3 Verification Study
As described in Section 4.2, the primary motivation behind the proposed variational prompt encoder is to generate more robust prompts that stay away from the boundaries of the target prompt region. To empirically validate whether we can truly generate prompts in the flatter regions of the loss landscape, we designed an experiment that does not require training a deep learning model, allowing for a direct comparison between the underlying principles of the variational and non-variational prompt encoders.
For the experiment of learning prompt embedding, we ran 200 experiments to generate 200 prompt embeddings. The initial prompt embeddings are randomly drawn from the normal distribution with a mean of 0 and a standard deviation of 25, and the loss function is formulated as the same loss function as VRP-SAM (see Equation 4). For the experiment of learning variational prompt distribution, we learn a single multivariate prompt distribution following the same formulation and reparameterization trick in Equation 5 and Equation 7 via gradient descent, given the loss function presented in Section 4.3. For both experiments, the gradient descent optimization process is stopped when the loss value does not improve by more than 0.0003 over 100 consecutive epochs. For other experimental settings not mentioned above, we follow the same practice as in Section 5.1.
@@ 5.4 Ablation Study
To thoroughly evaluate the effectiveness of different components of ProSAM, we conduct ablation studies on three different aspects: formulations of prompt distribution, inference strategies, and choices of image encoder. Due to the page limit, the quantitative comparison against VRP-SAM with the same number of learnable parameters as ours, the effects of different inference strategies, and more choices of image encoder, are presented in section 9.
Formulations of Prompt Distribution. As discussed in Section 4.2, a straightforward approach to modeling the prompt distribution $P(\bm{z}|I_{r},M_{r},I_{t})$ is to assume it follows a multivariate Gaussian distribution. Consequently, we perform an ablation study under this assumption, applying the reparameterization trick inspired by VAE [13]. As shown in Table 4, using the same number of Monte Carlo samples, ProSAM with Gaussian prompt distributions consistently underperforms compared to t-distributions across all folds of PASCAL-5i. This result highlights that the heavy-tailed nature of the t-distribution encourages a more robust mean prompt by increasing the likelihood of sampling outlier prompts, which in turn significantly enhances our testing performance on novel classes.
@@
Methods Parameters PASCAL-5i
K $\nu$ F-0 F-1 F-2 F-3 Means
ProSAM 10 5 75.26 77.57 70.29 65.22 72.04
ProSAM w/ Gaussian Dist 10 - 74.2 77.11 69.67 64.66 71.41
@@
Table 5: Ablation study on different image encoders, including DINOv2 [26].
Methods Image Encoder COCO-20i
@@ 7.2 The Robustness of ProSAM by Penalizing Laplacian
of the loss function $L$ with respect to the prompt embedding $z$, evaluated at the mean prompt $\mu_{z}$. The overall curvature is then quantified by the trace of the Hessian,
$\operatorname{Tr}\big{(}\nabla^{2}\mathcal{L}(\mu_{z})\big{)}=\sum_{i=1}^{n}\lambda_{i},$
where $\lambda_{i}$ are the eigenvalues of $\nabla^{2}\mathcal{L}(\mu_{z})$. Near a local minimum, where the segmentation loss is minimized, the loss function is typically convex or locally convex, ensuring that all eigenvalues satisfy $\lambda_{i}\geq 0$. Consequently, the Laplacian,
$\Delta\mathcal{L}(\mu_{z})=\operatorname{Tr}\big{(}\nabla^{2}\mathcal{L}(\mu_{z})\big{)},$
is nonnegative. When noise $\epsilon$ with variance $\sigma^{2}$ is added, a Taylor expansion of the loss function yields
$\mathbb{E}[\mathcal{L}(\mu_{z}+\epsilon)]=\mathcal{L}(\mu_{z})+\frac{1}{2}\sigma^{2}\,\Delta\mathcal{L}(\mu_{z})+O(\sigma^{3}).$
This extra term, $\frac{1}{2}\sigma^{2}\,\Delta\mathcal{L}(\mu_{z})$, penalizes high curvature regions, effectively encouraging the optimization process to favor flat minima. For ProSAM, this is crucial because a flat loss landscape implies that the predicted mean prompt $\mu_{z}$ is robust to small perturbations, thereby enhancing the stability and generalization of segmentation performance, particularly on novel objects.
@@ 7.3 The Limitation of VRP-SAM Without Laplacian Regularization
where $\mu_{z}$ is the learned prompt embedding, $f_{S}^{M}$ denotes the SAM mask decoder, $F_{S}^{I}(I_{t})$ is the image feature extraction, and $M_{t}$ represents the ground truth mask. This objective ensures that the generated mask is close to the target mask but does not explicitly encourage the embedding $\mu_{z}$ to reside in the center of the target prompt region $R_{I_{r},M_{r},I_{t}}$. As a result, the learned embedding can end up near the boundary of this region, where the loss function may exhibit high curvature.
Mathematically, the curvature at the embedding $\mu_{z}$ is characterized by the Hessian $\nabla^{2}\mathcal{L}(\mu_{z})$ of the loss function, and its trace, the Laplacian,
$\Delta\mathcal{L}(\mu_{z})=\operatorname{Tr}\big{(}\nabla^{2}\mathcal{L}(\mu_{z})\big{)},$
can be large if $\mu_{z}$ is near a boundary or a sharp region of the loss landscape. Without a regularization term that penalizes this curvature—such as the additional term $\frac{1}{2}\sigma^{2}\,\Delta\mathcal{L}(\mu_{z})$ obtained via noise injection—the optimizer is not explicitly guided to find flatter regions. Consequently, small perturbations in the embedding can lead to significant increases in loss, making the model more sensitive to noise and less robust. This sensitivity is particularly problematic when segmenting novel objects, where the embedding must generalize well to unseen variations. Hence, the absence of Laplacian regularization in VRP-SAM can result in unstable prompt embeddings and degraded segmentation performance.

### v3

@@ Abstract
The recent advancements in large foundation models have driven the success of open-set image segmentation, a task focused on segmenting objects beyond predefined categories. Among various prompt types (such as points, boxes, texts, and visual references), visual reference segmentation stands out for its unique flexibility and strong zero-shot capabilities. Recently, several SAM-based methods have made notable progress in this task by automatically generating prompts to guide SAM. However, these methods often generate prompts at boundaries of target regions due to suboptimal prompt encoder, which results in instability and reduced robustness. In this work, we introduce ProSAM, a simple but effective method to address the stability challenges we identified in existing SAM-based visual reference segmentation approaches. By learning a variational prompt encoder to predict multivariate prompt distributions, ProSAM avoids generating prompts that lie in unstable regions, overcoming the instability caused by less robust prompts. Our approach consistently surpasses state-of-the-art methods on the Pascal-5i and COCO-20i datasets, providing a more robust solution for visual reference segmentation.
@@ 1 Introduction
Figure 1: The same mask can be generated by SAM using various prompts in a region.
To tackle the challenges resulting from long-tailed data shortage and descriptive limitation, image segmentation with visual references (i.e., an annotated reference image that indicates the objects of interest) has become increasingly important. Its ultimate goal is to segment similar objects as indicated in the annotated reference image, regardless of the semantic category of target objects. Recently, various methods [45, 20, 32] have leveraged the exceptional segmentation capabilities of SAM, achieving significant breakthroughs in visual reference segmentation. The main idea of these methods is to generate prompts that direct SAM to predict masks for the target objects. Among these approaches, training-based methods [32] that learn prompt embeddings to guide SAM mask generation have achieved state-of-the-art (SOTA) performance. However, the existing training approaches often direct the prompts toward the boundaries of target regions, resulting in instability and reduced robustness. This robustness challenge is specifically caused by the inherent design of SAM — identical masks can be generated from different prompts (see Figure 1). This oversight poses a significant challenge for the training-based methods, where stability and generalizability are essential for zero-shot capability.
@@
  • We identify a commonly overlooked limitation in the SOTA SAM-based visual reference segmentation approach, where prompts are often generated at boundaries of target regions, leading to instability and reduced robustness.
@@ 3.2 Existing Training-Based Method Using SAM
In essence, VRP-SAM focuses solely on mask-level differences, while overlooking the potential for further optimizing the prompt encoder to generate more robust prompts.
@@ 4.1 Robustness Challenge
Figure 3: Intuitive illustration of the high-level idea behind the proposed variational prompt encoder. The dashed arrow shows the sampling and prompting procedure during training, while the solid arrow shows the prompting strategy during inference. Note that our generated prompts can be any type of prompt, while this illustration shows a positive point prompt as an example.
@@ 4.2 Variational Prompt Encoder
A straightforward instantiation of $P(\bm{z}|I_{r},M_{r},I_{t})$ is assuming $\bm{z}$ follows a conditional multivariate Gaussian distribution. However, considering our motivation is to push the mean prompt $\bm{\hat{\mu}}_{\bm{z}}$ toward the center of the target prompt embedding region, a more heavy-tailed t-distribution would be a better choice due to its statistical properties of having more chance to sample outliers [5]. In other words, when $\bm{\hat{\mu}}_{\bm{z}}$ is close to the boundary of $\mathcal{R}_{I_{r},M_{r},I_{t}}$, it is more likely to sample a prompt falling outside the target prompt embedding region (green region in Figure 3) if it follows heavy-tailed t-distribution. Given that the outlier prompt is more likely to result in low-quality mask predictions, the force to push the mean prompt toward the center of $\mathcal{R}_{I_{r},M_{r},I_{t}}$ will be greater with larger gradients. Theoretically, we also show in Section 7.5 that t-distribution results in a larger 4th-order curvature penalty that forces an extra push toward flatter and more stable regions. Inspired by Kim et al. [12], we formulate our variational prompt distribution $q_{\phi}(\bm{z}|I_{r},M_{r},I_{t})$ with the heavy-tailed property and diagonal covariance as follows,
@@ 5.1 Experimental Setup
Methods
Image
Encoder
Learnable
Params
COCO-20i PASCAL-5i
@@ F-0 F-1 F-2 F-3 Mean F-0 F-1 F-2 F-3 Mean
Implementation Details. To ensure a strictly fair comparison with VRP-SAM, the SOTA method in the visual reference segmentation, we ensure that all the experimental settings (e.g., random seed, LR scheduler, optimizer) and hyper-parameters (e.g., number of layers, prompt embedding dimensions, number of prompts) are identical to VRP-SAM. Specifically, the model architecture of the variational prompt encoder is similar to VRP-SAM (see Section 8 in Appendix for more details), except two linear layers have been added at the end of the prompt encoder to predict the mean and log standard deviation of the multivariate prompt distribution. In other words, our method is easy to implement with only a few lines of code, but can effectively boost the robustness of generated prompts. During training, the number of Monte Carlo samples K has been set to 10, with the degrees of freedom $\nu$ of 5. Also, we employ the AdamW [21] optimizer with weight decay of 1e-6 and the cosine annealing learning rate scheduler with warm restart [22] after 15 epochs, and the initial learning rate of 1e-4. The model is trained with 100 epochs with a fixed random seed of 321. For the choice of image encoder utilized by the variational prompt encoder, ResNet-50 [7] is adopted following VRP-SAM. In addition, our variational prompt encoder predicts 50 multivariate prompt distributions per target image and only uses 50 mean prompts during inference, which guarantees the same number of prompts have been used in the experimental study for both VRP-SAM and ProSAM. In terms of the visual annotation types, we follow the same procedure as VRP-SAM [32] and SEEM [46] to automatically generate points, scribbles, and boxes based on the mask annotations. Lastly, the mean intersection over union (mIOU) is adopted to evaluate our segmentation performance across all datasets. All experiments on COCO-20i were conducted on 4 RTX 4090 GPUs with a batch size of 2 per GPU, whereas the experiments on Pascal-5i were conducted on 1 H100 GPU with a batch size of 2. On a single H100 GPU with batch size 2, ProSAM takes 0.19s per training batch and 0.12s per inference batch.
@@ 5.3 Verification Study
As described in Section 4.2, the primary motivation behind the proposed variational prompt encoder is to generate more robust prompts that lie away from the boundaries of the target prompt region $\mathcal{R}_{I_{r},M_{r},I_{t}}$. To empirically validate this, we conducted two complementary studies to assess the robustness of prompts generated by ProSAM. Due to the page limit, additional verification studies are included in Appendix Section 9.
Robustness to Noise Perturbation. If our generated prompts truly stay away from the boundaries of the target prompt region, then a small perturbation applied in the latent prompt embedding space should result in minimal degradation in mask quality. To evaluate this, we injected Gaussian noise into the prompt embeddings of both ProSAM and VRP-SAM during inference and analyzed their robustness. As shown in Figure 4, when the standard deviation of Gaussian noise is set to 1.2, the mIoU of VRP-SAM degrades by approximately 40%, while ProSAM only shows a 20% drop. Overall, VRP-SAM exhibits a much greater sensitivity to noise, suggesting that its prompts tend to lie near the boundaries of target prompt regions, whereas ProSAM prompts remain significantly more stable under perturbations, providing evidence that they are located closer to the center of the target prompt region.
Figure 4: Gaussian noise perturbation on ProSAM prompts and VRP-SAM prompts in the latent space. Both ProSAM model and the VRP-SAM model are trained on PASCAL-5i F-0, and the noise perturbation is injected during inference.
Proximity to Center Prompts. We further assess prompt robustness by measuring the similarity between the predicted prompts and center point prompts in the ground-truth masks. Since the center prompt of the target prompt regions is unobservable, we approximate it using the latent prompt embedding of the center point prompt in the ground-truth masks because those are the most robust prompts we can obtain. Specifically, we identified the center pixels in the ground-truth masks and fed them into the SAM prompt encoder to get their corresponding prompt embeddings. Then, for each fold in PASCAL-5i, we computed the cosine similarity between the predicted prompts and the embeddings of the center point prompts. As shown in Table 4, ProSAM prompts consistently exhibit higher cosine similarity to the center point prompts than VRP-SAM prompts. This observation further supports our hypothesis that ProSAM generates prompts more aligned with the central and robust regions of the target prompt region. Additionally, we evaluated the statistical relationship between cosine similarity and segmentation quality by computing the Pearson correlation coefficient between cosine similarity and mIoU. A positive correlation coefficient of 0.365 suggests that prompts closer to the center point prompt are more likely to yield high-quality segmentation results.
Table 4: The average cosine similarity between predicted prompts and the embeddings of center point prompts in ground-truth mask.
Cosine Similarity
to Center Prompts
Methods
VRP-SAM ProSAM
F-0 0.0072 0.0122
F-1 0.0127 0.0291
F-2 -0.0040 0.0341
F-3 0.0106 0.0420
@@ 5.4 Ablation Study
To thoroughly evaluate the effectiveness of different components of ProSAM, we conduct ablation studies on three different aspects: formulations of prompt distribution, inference strategies, and choices of image encoder. Due to the page limit, our comparison against VRP-SAM with the same number of learnable parameters and more choices of image encoder, is presented in Appendix Section 10 .
Formulations of Prompt Distribution. As discussed in Section 4.2, a straightforward approach to modeling the prompt distribution $P(\bm{z}|I_{r},M_{r},I_{t})$ is to assume it follows a multivariate Gaussian distribution. Consequently, we perform an ablation study under this assumption, applying the reparameterization trick inspired by VAE [13]. As shown in Table 5, with different $K$ and $\nu$, ProSAM with Gaussian prompt distributions consistently under-perform compared to t-distributions. This result highlights that the heavy-tailed nature of the t-distribution encourages a more robust mean prompt by enforcing a larger 4th-order curvature penalty, which in turn significantly enhances our testing performance on novel classes.
@@
Method Parameters PASCAL-5i Method Parameters PASCAL-5i
K $\nu$ Means K $\nu$ Means
ProSAM 10 5 72.04
ProSAM
w/ Gaussian
10 - 71.41
10 3 71.55
15 5 71.42 15 - 71.01
15 3 71.02
@@
Table 6: Ablation study on different image encoders.
Methods
Image
Encoder
COCO-20i
@@ F-0 F-1 F-2 F-3 Mean
Inference Strategies. During inference, the predicted mean prompt $\bm{\hat{\mu}}_{\bm{z}}$ is employed to generate the predicted mask, since the robustness of $\bm{\hat{\mu}}_{\bm{z}}$ can be guaranteed due to the existence of margin between $\bm{\hat{\mu}}_{\bm{z}}$ and the boundary of target prompt region $\mathcal{R}_{I_{r},M_{r},I_{t}}$. However, it is still interesting to see the effectiveness of utilizing randomly sampled prompts from the learned prompt distribution during inference. Therefore, we experimented with 5 different inference strategies, in which we sample $K$ prompts and merge their corresponding $K$ masks in 5 different ways. Specifically, they merge either the logit masks or the binary masks given a threshold of 0.5. From Table 7, all the inference strategies with different ways of merging $K$ masks achieve comparable results to using only the mask prompted by the $\bm{\hat{\mu}}_{\bm{z}}$. This indicates that our learned multivariate prompt distribution spans over the target prompt region.
Methods Inference Strategies mIOU
ProSAM mean prompt only 75.26
ProSAM
max of $K$ logit masks
74.69
mean of $K$ logit masks
75.16
max of $K$ binary masks
74.67
mean of $K$ binary masks
74.92
majority vote of $K$ binary masks
75.05
@@ 7.2 The Robustness of ProSAM by Penalizing Laplacian
of the loss function $L$ with respect to the prompt embedding $z$, evaluated at the mean prompt $\mu_{z}$. The overall curvature is then quantified by the trace of the Hessian, namely the Laplacian,
$\Delta\mathcal{L}(\mu_{z})=\operatorname{Tr}\big{(}\nabla^{2}\mathcal{L}(\mu_{z})\big{)}=\sum_{i=1}^{n}\lambda_{i},$
where $\lambda_{i}$ are the eigenvalues of $\nabla^{2}\mathcal{L}(\mu_{z})$. Near a local minimum, where the segmentation loss is minimized, the loss function is typically convex or locally convex, ensuring that all eigenvalues satisfy $\lambda_{i}\geq 0$. Consequently, the Laplacian $\Delta\mathcal{L}(\mu_{z})$ is nonnegative. Under this constraint, minimizing the Laplacian strictly leads to lower overall curvature. Following the conclusion from Section 7.1, when noise $\epsilon$ with variance $\sigma^{2}$ is added, the Laplacian is implicitly penalized, which effectively encourages the optimization process to favor flat minima over high curvature regions. For ProSAM, this is crucial because a flat loss landscape implies that the predicted mean prompt $\mu_{z}$ is robust to small perturbations, thereby enhancing the stability and generalization of segmentation performance, particularly on novel objects.
@@ 7.3 The Limitation of VRP-SAM Without Laplacian Regularization
where $\mu_{z}$ is the learned prompt embedding, $f_{S}^{M}$ denotes the SAM mask decoder, $F_{S}^{I}(I_{t})$ is the image feature extraction, and $M_{t}$ represents the ground truth mask. This objective ensures that the generated mask is close to the target mask but does not explicitly encourage the embedding $\mu_{z}$ to reside in the low-curvature area of the target prompt region $R_{I_{r},M_{r},I_{t}}$. As a result, the learned embedding may end up in an area where the loss function exhibits high curvature.
As explained in Section 7.2, the curvature at the embedding $\mu_{z}$ is characterized by the Hessian $\nabla^{2}\mathcal{L}(\mu_{z})$ of the loss function, and its trace, the Laplacian $\Delta\mathcal{L}(\mu_{z})$ can be large if $\mu_{z}$ is near a boundary or a sharp region of the loss landscape. Without a regularization term that penalizes this curvature—such as the additional term $\frac{1}{2}\sigma^{2}\,\Delta\mathcal{L}(\mu_{z})$ obtained via noise injection—the optimizer is not explicitly guided to find flatter regions. Consequently, small perturbations in the embedding can lead to significant increases in loss, making the model more sensitive to noise and less robust. This sensitivity is particularly problematic when segmenting novel objects, where the embedding must generalize well to unseen variations. Hence, the absence of Laplacian regularization in VRP-SAM can result in unstable prompt embeddings and degraded segmentation performance.